Find the centre and radius of the equation of the circle. 2x^2+2y^2-8x+5y+10=0
1
Expert's answer
2020-12-10T08:48:28-0500

Explanations & Calculations


  • A rearrangement is needed to identify the needed quantities.

"\qquad\qquad
\begin{aligned}
\small 2x^2 +2y^2-8x+5y+10&= \small0\\
\small x^2 +y^2 -4x+\frac{5}{2}y+5&=\small 0\\
\small \big[x^2-4x\big]+\big[y^2+\frac{5}{2}y\big]+5&=\small 0\\
\small \big[(x-2)^2-4\big]+\big[(y+\frac{5}{4})^2-\frac{25}{16}\big]+5&=\small 0\\
\small (x-2)^2+(y+\frac{5}{4})^2-\frac{9}{16}&= \small 0\\
\small \Big[x-2\Big]^2+\Big[y-(\frac{5}{4})\Big]^2 &=\small \Big(\frac{3}{4}\Big)^2\\

\end{aligned}"


  • By comparing this equation with that used at the derivation of the equation of a circle "\small (x-f)^2+(y-g)^2=r^2" ,


  • Therefore,
  • Centre = "(2,-\frac{5}{4})"
  • Radius = "\frac{3}{4}"